OpenStudy (beth12345):

how to find x when x/3 + (1/(x-2)) =2

OpenStudy (beth12345):

\[\frac{ x }{ 3 } + \frac{ 1 }{ x-2 } = 2\] i find it easier to understand when its written in this form

OpenStudy (campbell_st):

multiply each term by 3(x-2) this will eliminate the fraction which will give \[x(x-2) + 3 = 2 \times 3(x-2)\] next you need to expand and simplify. Write the equation will all terms on the left and equal to zero. Then you need to solve the resulting quadratic

OpenStudy (beth12345):

how do i get both sides to be multiplied by x-2?

OpenStudy (beth12345):

like wouldn't it cancel out?

OpenStudy (campbell_st):

(x - 2) is the denominator of the 2nd term. the common denominator is 3(x-2) by multiplying by the common denominator you eliminate all denominators

OpenStudy (beth12345):

ok

OpenStudy (beth12345):

so would it become x+1=2X3(x-2) ?

OpenStudy (beth12345):

\[x+1=2\times 3(x-2)\]

OpenStudy (beth12345):

oh ok nvm, i see how you did that

OpenStudy (beth12345):

i was multiplying everything on the left as one term

OpenStudy (beth12345):

would it afterward become (x-5)(x-3) ?

OpenStudy (beth12345):

=0 *

OpenStudy (beth12345):

and then x= 5 or 3 and x=/=2 ?

OpenStudy (beth12345):

ok ya i got it, thanks